# NLP-enabled trajectory map-matching in urban road networks using transformer sequence-to-sequence model

### v1

@@ V-D Implementation Detail
After evaluating different sets, model hyperparameters are configured as follows. The embedding dimension ($d_{emb}$) is set to 512, while the MLP intermediate dimension ($d_{ff}$) is set to 2048. We use 8 transformer blocks for both the encoder and decoder, and each multi-head attention block employs 16 linear transformations, resulting in $d_{k}=32$. During training, we utilize the Adam optimizer with a learning rate of 0.0001 and the cross-entropy loss function. The batch size is fixed at 200, and the model undergoes training for a total of 32 epochs. For the RNN model, the embedding dimension is set to 1024. We employ the Adam optimizer with a learning rate of 0.0005 for both the encoder and decoder. Similarly, the batch size is fixed at 200, and the model undergoes training for a total of x epochs reaching to the point that the loss function converges. We implemented the model in Python using the PyTorch library. We trained our models on one machine with NVIDIA RTX 6000 GPUs.
V-E Evaluation Metrics
The performance of the model is evaluated using four different metrics, accuracy, length-based accuracy, Jackkard similarity, and Bilingual Evaluation Understudy (BLEU) [58], mathematically described below:
$\text{Accuracy}=\textit{Acc}_{1}=\frac{1}{N}\sum_{i=1}^{N}\frac{|\hat{T}_{i}\cap T_{i}|}{|T_{i}|}$
[…]
where $T_{i}$ and $\hat{T}_{i}$ are the $i$th ground truth path and the estimated path in the whole trajectory set $T$. |.||.|| . | represents the segment count of the trajectory while $lenght(.)$ represents the length of the trajectory in meters.
The Jaccard similarity is a measure of similarity between two sets. Here, we use this metric to compare each trajectory with model estimation. It is defined as the size of the intersection of the two routes composed of road segments divided by the size of their union and yields a value between 0 and 1, where 0 indicates no similarity (no common elements) and 1 indicates complete similarity (identical sets).
$\text{Jackkard}=J(T_{i},\hat{T}_{i})=\frac{1}{N}\sum_{i=1}^{N}\frac{|\hat{T}_{i}\cap T_{i}|}{|\hat{T}_{i}\cup T_{i}|}$
BLEU is a metric commonly used to evaluate the quality of machine-generated text, particularly in the context of machine translation. However, in the context of this paper, BLEU is utilized to evaluate the similarity of the predicted segment-based trajectory with the ground truth trajectory. Similar to the text evaluation, here it is assumed to operate by comparing n-grams (contiguous sequences of n road segments) in the candidate and reference trajectories. The intuition behind BLEU is that a good mapping should have similar n-gram statistics to those of the ground truth trajectories. It has two parts, precision, and brevity penalty. Precision which counts how many contiguous sequences of “n” items (n-grams) in the candidate route match those in the reference route and divides this count by the total number of n-grams in the candidate routes and brevity penalty that penalizes the short mappings [58].
$Precision(\hat{T}_{i},T_{i})=\frac{\text{count of n-gram matches between }\hat{T}_{i},T_{i}}{\text{Count of n-grams in }\hat{T}_{i}},$
$BP=\begin{cases}1&|\hat{T}_{i}|>|T_{i}|\\ \exp{(1-\frac{|\hat{T}_{i}|}{|T_{i}|})}&|\hat{T}_{i}|\leq|T_{i}|,\\ \end{cases},$
$BLEU=BP\exp(\frac{1}{n}\sum_{i=1}^{n}\log(Precision_{i}))$
VI Results
@@
The trajectory data is randomly split into 98% (31,455 trajectories) for training and 2% (642 trajectories) for testing. The average length of trajectories used for testing the model’s performance is 16 kilometers (10 miles). After trajectory segmentation to adhere to the constraint of the maximum length, the data size expands to 272,722 to train the transformer model and 402,361 to train the Bi-GRU encoder-decoder model. The models are trained using the train data set and then evaluated with the test data set. The transformer encoder-decoder model outperformed the Bi-GRU encoder-decoder and resulted in an accuracy ($\textit{Acc}_{1}$) of 75.2%, length-based accuracy ($\textit{Acc}_{2}$) of 76.1%, Jackkard similarity (J) of 78.1%, and BLEU of 75.6%. While, the Bi-GRU encoder-decoder model resulted in the accuracy ($\textit{Acc}_{1}$) of 58.3%, length-based accuracy ($\textit{Acc}_{2}$) of 58.5%, Jackkard similarity (J) of 53.7% and BLEU of 53.9%. Results are summarized in table I. The transformer model reached convergence after 30 epochs, with each epoch completing in 33 minutes. On the other hand, the RNN-based model achieved convergence in 10 epochs, although each epoch took a longer 1632 minutes. However, a closer examination of the inference times for various trajectory lengths, as depicted in Figure 7, reveals that the RNN-based encoder-decoder model exhibits faster inference times than the transformer model. This variance can be attributed to the distinctive architectural characteristics of the models. Generally, the transformers-based encoder-decoder model exhibits more effective and faster convergence compared to the RNN-based counterpart. Several factors contribute to this phenomenon. First, the self-attention layer’s sequential operation complexity is $O(1)$ while the recurrent layers’ sequential operation is $O(n)$. The self-attention layer captures relationships and dependencies between different parts of the input sequence. Then, the number of operations required to connect all positions (or elements) within the sequence remains constant, regardless of the sequence length. This is typically achieved by using matrix operations that can be efficiently paralleled, making self-attention well-suited for capturing long-range dependencies in sequences. A recurrent layer, on the other hand, involves operations that depend on the previous states of the layer. Thus, the computation at each step is dependent on the computation of the previous step. As a result, the number of sequential operations required by a recurrent layer scales linearly with the length of the sequence. This is represented by $O(n)$, where $n$ is the length of the sequence. Second, the complexity per layer of the self-attention layer used in the transformer is $O(n^{2}d)$ where $d$ is the model dimension (in our case the model embedding dimension). In a self-attention layer, attention scores are computed between all pairs of positions in the input sequence, resulting in a quadratic complexity with respect to the sequence length $n$. Additionally, each attention score involves a dot-product computation between two $d$-dimensional vectors, which contributes to the $d$ factor in the complexity. While the recurrent layer’s complexity is $O(nd^{2})$. At each position, there’s a multiplication of a $d$ dimensional vector (the input) with a $d$ dimensional matrix (the recurrent weight matrix), leading to the $d^{2}$ factor in the complexity expression and the overall complexity of $O(nd^{2})$ Since in our study, the model dimension is larger than the input sequence, thus, the recurrent layer is more complex than the self-attention layer.
@@
Model $\textit{Acc}_{1}$ $\textit{Acc}_{2}$ $J$ BLEU
Transformer 0.752 0.761 0.781 0.756
RNN(BiGRU) 0.583 0.585 0.537 0.539
TABLE I: Models’ Performance Metrics

### v2

@@ VI-D Implementation Detail
After evaluating different sets, model hyperparameters are configured as follows. The embedding dimension ($d_{emb}$) is set to 512, while the MLP intermediate dimension ($d_{ff}$) is set to 2048. We use 8 transformer blocks for both the encoder and decoder, and each multi-head attention block employs 16 linear transformations, resulting in $d_{k}=32$. During training, we utilize the Adam optimizer with a learning rate of 0.0001 and the cross-entropy loss function. The batch size is fixed at 200, and the model undergoes training for a total of 32 epochs. For the RNN model, the embedding dimension is set to 1024. We employ the Adam optimizer with a learning rate of 0.0005 for both the encoder and decoder. Similarly, the batch size is fixed at 200, and the model undergoes training for a total of 32 epochs reaching to the point that the loss function converges. We implemented the model in Python using the PyTorch library and trained it on a machine equipped with NVIDIA RTX 6000 GPUs.
VII Results
@@
The trajectory data is randomly split into 98% (31,455 trajectories) for training and 2% (642 trajectories) for testing. The complexity of the Manhattan road network with around 17,965 unique segments, posed a significant challenge given our relatively sparse dataset of 31,455 trajectories. To ensure comprehensive coverage of the road network and effective learning from this data, we opted for a split of 98% for train and 2% for test (resulting in 642 unseen trajectories for testing the model). This decision was made to maximize the model’s exposure to the data covering the complex road network of Manhattan, which was crucial for training. However, despite this high training proportion, our model demonstrated robust performance on the unseen data, indicating that it did not overfit and retained its generalization capability. The average length of trajectories used for testing the model’s performance is 16 kilometers (10 miles). After trajectory segmentation to adhere to the constraint of the maximum length, the data size expands to 272,722 to train the transformer model and 402,361 to train the Bi-GRU encoder-decoder model. The models are trained using the train data set and then evaluated with the test data set. The transformer encoder-decoder model outperformed the Bi-GRU encoder-decoder and resulted in an accuracy of 75.2%, Jackkard similarity (J) of 78.1%, and BLEU of 75.6%. While, the Bi-GRU encoder-decoder model resulted in the accuracy of 58.3%, Jackkard similarity (J) of 53.7% and BLEU of 53.9%. Results are summarized in table II. The transformer model reached convergence after 32 epochs, with each epoch completing in 33 minutes. On the other hand, the RNN-based model achieved convergence in 10 epochs, although each epoch took a longer 1632 minutes. However, a closer examination of the inference times for various trajectory lengths, as depicted in Figure 8, reveals that the RNN-based encoder-decoder model exhibits faster inference times than the constructed transformer model. This variance can be attributed to the distinctive architectural characteristics of the models. Generally, the transformers-based encoder-decoder model exhibits more effective and faster convergence compared to the RNN-based counterpart. Several factors contribute to this phenomenon. First, the self-attention layer’s sequential operation complexity is $O(1)$ while the recurrent layers’ sequential operation is $O(n)$. The self-attention layer captures relationships and dependencies between different parts of the input sequence. Then, the number of operations required to connect all positions (or elements) within the sequence remains constant, regardless of the sequence length. This is typically achieved by using matrix operations that can be efficiently paralleled, making self-attention well-suited for capturing long-range dependencies in sequences. A recurrent layer, on the other hand, involves operations that depend on the previous states of the layer. Thus, the computation at each step is dependent on the computation of the previous step. As a result, the number of sequential operations required by a recurrent layer scales linearly with the length of the sequence. This is represented by $O(n)$, where $n$ is the length of the sequence. Second, the complexity per layer of the self-attention layer used in the transformer is $O(n^{2}d)$ where $d$ is the model dimension (in our case the model embedding dimension). In a self-attention layer, attention scores are computed between all pairs of positions in the input sequence, resulting in a quadratic complexity with respect to the sequence length $n$. Additionally, each attention score involves a dot-product computation between two $d$-dimensional vectors, which contributes to the $d$ factor in the complexity. While the recurrent layer’s complexity is $O(nd^{2})$. At each position, there’s a multiplication of a $d$ dimensional vector (the input) with a $d$ dimensional matrix (the recurrent weight matrix), leading to the $d^{2}$ factor in the complexity expression and the overall complexity of $O(nd^{2})$ Since in our study, the model dimension is larger than the input sequence, thus, the recurrent layer is more complex than the self-attention layer.
@@
TABLE II: Models’ Performance Metrics
Model Accuracy J BLEU
Transformer 0.752 0.781 0.756
RNN(BiGRU) 0.583 0.537 0.539